OpenStudy (anonymous):

What is the solution to the system of equations? 2x-3y-2z=-7 x-3y-3z=4 x-2y-z=-3

OpenStudy (zzr0ck3r):

x=2y+z-3 so (2y+z-3)-3y-3z=4 -y-2z-3=4 so y = -7-2z so x = 2(-7-2z)+z-3 so 2x-3y-2z=-7 ==>2(2(-7-2z)+z-3)-3(-7-2z)-2z=-7 -28-8z+z-3+21+6z-2z=-7 -8z+z+6z=3 z=-3 y = -1 x = -8

OpenStudy (zzr0ck3r):

make sense @DribleDaVaca19 ?